0.2 ml of a 10-9 dilution of a bacterial suspension plated on an agar plate gave rise to 156 distinct colonies. What is the concentration (CFU/ml) of bacteria in the original, undiluted suspension?
Answer #1

CFU/ml or colony forming units/ ml = (number of colonies X dilution factor)/ volume of suspension plated

Number of colonies = 156

Dilution factor = 1/ dilution = 1/10-9 = 109

Volume of bacterial suspension plated on agar = 0.2ml

CFU/ml) of bacteria in the original = (156 X 109)/0.2 = 780 X 109 = 7.8 X 1011 CFU / ml in the original suspension

So the original undiluted suspension had 7.8 X 1011 Colony forming units per ml